what are the angles of a triangle that has sides of 21 28 and 30
Answer by Alan3354(30983)

Use the law of cosines:
c^2 = a^2 + b^2 - 2abcos(C)
900 = 441 + 784 - 2*21*28*cos(C)
cos(C) =~ 0.27636
C =~ 73.96º
---------------
Do the other 2 the same way, or use the Law of Sines